OpenStudy (anonymous):

Determine if the function is a polynomial function. if it is, identify the degree g(X)=xexponent 1over 5 + pie x exponent3 - two over five x

OpenStudy (anonymous):

i this is not a polynomial function...

OpenStudy (anonymous):

g(x)=x/5+pix^3-(2/5)x? or g(x)=x/5+pix^3-2/(5x)?

OpenStudy (anonymous):

this is because the last term u hav 2/5x which gives u a negative index and so it no longer remains a polynomial :)

OpenStudy (anonymous):

o ok thank you